OpenStudy (anonymous):

Find the value of the variable(s). If your answer is an integer, leave it in simplest radical form. The triangle is not drawn to scale.

OpenStudy (anonymous):

\[a) x=10\sqrt{3}, y=30\] \[b) x=10, y=30\sqrt{3}\] \[c) x=30\sqrt{3}, y=10\] \[d) x=30, y=10\sqrt{3}\]

OpenStudy (nikato):

It's a 30-60-90 triangle so the sides form a special ratio http://www.regentsprep.org/Regents/math/algtrig/ATT2/Ltri30.htm To help solve. You can look at short cut pattern formulas. If you look at your diagram, you should see that you were given the long leg(LL)

OpenStudy (anonymous):

idk if I did this problem right , but i think it's D.

OpenStudy (whpalmer4):

Yes, that should be correct.

OpenStudy (nikato):

Yes D is correct